Sara has $40 in her budget. Given the following graph, what is the MRS of her optimal consumption bundle? Burritos Pizza 4
Answer #1

Ans. Income, M = $40

Price of Pizza = Px

Price of Burritos = Py

Thus, y- intercept which represents maximum consumption of burritos = 8 = M/Py

=> Py = $5

and x-intercept which represents maximum consumption of pizza = 4 = M/Px

=> Px = $10

At optimal consumption bundle,

MRS = -Px/Py (Slope of budget constraint)

=> MRS = -10/5 = -2

Thus, MRS = -2

Thank you